# A Unified Framework for Direction and Diffuseness Estimation Using Tight-Frame Microphone Arrays

### v2

@@ VI Conclusion
This study presented a comparative investigation of diffuseness estimation methods based on intensity-energy relations and eigenvalue analysis, using several microphone arrays including the Afmt, Fibo64, and TF24 systems. Through a series of controlled simulations, the relationships between the diffuseness index and the beam-to-diffuse ratio were systematically examined under single-source, mixed, and interference conditions. The results demonstrated that the eigenvalue-based indices, particularly $\Psi_{\mathrm{COM}}$, exhibited higher robustness and a more linear response than the conventional intensity-energy (I/E) approach. The I/E method tended to overestimate diffuseness in interference fields where the active intensity approaches zero, leading to physically inconsistent interpretations.
Regarding the combination of microphone array configurations and evaluation approaches, the Fibo64 array would be the first choice if such an array is available. In cases where this is not realistic, arrays composed of microphones with known directivity patterns—such as the Tight frame configuration, which possesses theoretical validity—represent a reasonable alternative. Furthermore, the eigenvalue analysis based on the covariance matrix constructed only from the FOA velocity components ($u_{x}$, $u_{y}$, and $u_{z}$) was found to produce stable and reliable results overall.

### v3

@@ VI Conclusion
This study presented a comparative investigation of diffuseness estimation methods based on intensity-energy relations and eigenvalue analysis, using several microphone arrays including the Afmt, Fibo64, and TF24 systems. Through a series of controlled simulations, the relationships between the diffuseness index and the beam-to-diffuse ratio were systematically examined under single-source, mixed, and interference conditions. The results demonstrated that the eigenvalue-based indices, particularly $\Psi_{\mathrm{COM}}$, exhibited higher robustness and a more linear response than the conventional intensity-energy (I/E) approach. The I/E method tended to overestimate diffuseness in interference fields where the active intensity approaches zero, which may lead to physically inconsistent interpretations.
Regarding the combination of microphone array configurations and evaluation approaches, the Fibo64 array would be the first choice if such an array is available. In cases where this is not realistic, arrays composed of microphones with known directivity patterns—such as the Tight frame configuration, which possesses theoretical well-defined structure—represent a reasonable alternative. Furthermore, the eigenvalue analysis based on the covariance matrix constructed only from the FOA velocity components ($u_{x}$, $u_{y}$, and $u_{z}$) was found to produce stable and reliable results overall.
@@
While the TF24 formulation does not aim at a strict physical equivalence to pressure-velocity measurements, the present study demonstrates that, once its assumptions and limitations are explicitly stated, the proposed pseudo-intensity-based formulation provides a practically useful and internally consistent basis for comparative diffuseness evaluation. In this sense, the TF24 framework should be interpreted not as a replacement for physically exact intensity measurements, but as an operational representation that enables stable and interpretable diffuseness estimation across different array geometries.